OpenStudy (anonymous):

Factor completely: 2x3 + 14x2 + 4x + 28

OpenStudy (anonymous):

I am stuck between two answers

OpenStudy (anonymous):

2(x3 + 7x2 + 2x + 14) (2x + 14)(x2 + 2)

pooja195 (pooja195):

2(x+7)(x^2+2)

pooja195 (pooja195):

i think?

OpenStudy (zehanz):

Split up in two times two terms: (2x³ + 14x²) + (4x + 28)=2x²(x+7) + 4(x+7)

OpenStudy (zehanz):

Now you can see that there is a "new" common factor: x+7. So it can be factored out: (2x²+4)(x+7)

OpenStudy (anonymous):

I agree with Pooja....2(x + 7)(x^2 + 2)

OpenStudy (anonymous):

so it would be the second choice?

OpenStudy (anonymous):

or would it be 2[(x + 7)(x2 + 2)]

OpenStudy (zehanz):

Of course, now the hard work has been doe, it's time to clean up: (2x²+4)(x+7)=2(x²+2)(x+7)

OpenStudy (anonymous):

thank you!

OpenStudy (zehanz):

So there really is no difference between @pooja195's answer and mine...

OpenStudy (zehanz):

YW!

pooja195 (pooja195):

:)

OpenStudy (zehanz):

What I don't get is this: you're supposed to factor *completely*, but you can choose between: 2(x3 + 7x2 + 2x + 14) and (2x + 14)(x2 + 2) both of which are not factored completely...